# A unique Neutrino Mass Matrix Texture under Exponential Parametrization

### v2

@@ Abstract
Under the exponential parametrization scheme $m^{\nu}_{ij}\sim re^{i\theta}$, we propose a unique neutrino mass matrix texture that exhibits four correlations among its elements. The mixing scheme obtained from the proposed texture is consistent with experimental observations. We construct a concrete model based on the $SU(2)_{L}\times U(1)_{Y}\times A_{4}\times Z_{10}$ group within the framework of the seesaw mechanism.
journal: arXiv
@@ 1 Introduction
The neutrino oscillation suggests that neutrinos change their identity as they traverse space [1]. This phenomenon can be understood from the quantum mechanics framework which tells us that the three flavour states of neutrinos ($\nu_{l=e,\mu,\tau}$) are the mixture of three mass eigenstates $(\nu_{i=1,2,3})$ having definite masses ($m_{i=1,2,3}$). The neutrino oscillation predicts the existence of non-zero mass, which goes against the predictions of the Standard Model (SM) [2, 3, 4]. In this regard, we need to go beyond the SM framework to understand the non-zero neutrino masses. The neutrino mass matrix ($M_{\nu}$) originates from the Yukawa Lagrangian, which can be understood from the see-saw mechanism [5, 6, 7, 8, 9, 10, 11, 12, 13]. In general, a Majorana neutrino mass matrix contains the information of twelve parameters: three mass eigenvalues ($m_{i=1,2,3}$), three mixing angles ($\theta_{12},\theta_{13},\theta_{23}$), three CP phases ($\delta,\alpha,\beta$), and three unphysical phases ($\phi_{1},\phi_{2},\phi_{3}$). Recent experiments predict three mixing angles and two mass-squared differences ($\Delta\,m_{21}^{2}$ and $\Delta\,m_{31}^{2}$) precisely [14]. However, the octant of $\theta_{23}$, neutrino mass ordering and a precise bound of $\delta$ are yet to be determined. In addition, the oscillation experiments do not measure three individual mass eigenvalues and the Majorana phases.
@@
However, the $M_{\nu}$ carries the information of all nine physical parameters, and by imposing certain constraints on $M_{\nu}$, some of the physical parameters can be predicted. Various constraints, such as texture zeroes [15], vanishing minors [16], hybrid textures [17], and $\mu-\tau$ symmetry [18] are widely studied in the literature. The experiments rule out the $\mu-\tau$ symmetry due to its prediction of $\theta_{13}$ being zero. In this regard, there are several attempts made to explore deviations from the exact $\mu-\tau$ symmetry [19, 20, 21]. In the present work, we explore different types of constraints, deviating from the existing literature. We adhere to the exponential parametrization of the neutrino mass matrix $M_{\nu_{ij}}=r_{ij}e^{i\theta_{ij}}$, where, $r_{ij}$ represents the absolute value and $\theta$ stands for argument of the matrix elements. In contrast, a complex number $z$ can also be expressed as $z=x+iy$. In principle, physics is independent of parametrization. The underlying principles and phenomena do not alter regardless of the parametrization we choose. However, a suitable parametrization may lead to understand the underlying physics in a better way. Like several other constraints studied in the literature, the said parametrization scheme offers new insights into how physical parameters can be correlated in a different way. The idea of the said parametrization was first discussed in a model-independent way in Ref [22]. The latter discusses several promising textures from different categories, highlighting distinct correlations including the exponential parametrization. All the proposed textures are experimentally viable. However, we have not provided any concrete model in support of the said textures.
@@ 2.1 The Proposed Texture
To draw information on the physical parameters, we need to diagonalize $M:M^{d}=V^{T}.M.V$, where, $V$ is a $3\times 3$ unitary matrix known as Pontecorvo-Maki-Nakagawa-Sakata (PMNS) matrix [23]. In general, the PMNS matrix is defined in flavour basis i.e., the basis where the charged lepton mass matrix is diagonal. The Particle Data Group (PDG) [24] has adopted a parametrization for $V$, where, $V$ is parametrized using three angles and six phases as shown below,
@@ 2.2 Numerical Analysis
The four correlations appearing in $M_{\nu}$ of Eq. (1), are manifested as four transcendental equations involving nine physical parameters: $f_{i}(\theta_{12},\theta_{13},\theta_{23},m_{1},m_{2},m_{3},\delta,\alpha,\beta)=0$, where, $i=1,2.3,4$. We solve these equations by the Newton-Raphson method to generate a sufficient number of random numbers for the parameters $\alpha$, $\beta$, $\delta$ and $\theta_{23}$. For this, we set the mixing angles $\theta_{12}$ and $\theta_{13}$ at their respective $3\sigma$ bound. The mass eigenvalues $m_{1}$, $m_{2}$ and $m_{3}$ are estimated from two mass squared differences $\Delta m_{21}^{2}$ and $\Delta m_{31}^{2}$, such that $\Sigma m_{i}$ is consistent with the cosmological data [25].
The analysis leads to some interesting predictions of the observable parameters. We see that $\theta_{23}$ lies within $42^{\circ}$ to $47^{\circ}$ approximately (within the $3\sigma$ range [14]). The $\delta$ experiences two narrow bounds: $5.43^{\circ}$-$12.32^{\circ}$; and $347.52^{\circ}$-$354.52^{\circ}$. Thus, we see that within the allowed $3\sigma$ bound for $\delta$ given by the experiments, the above texture defines a forbidden region also. Similarly, we see that the Majorana phases $\alpha$ and $\beta$ have got two and four allowed bounds respectively. We see, $\alpha\sim 85.37^{\circ}-94.66^{\circ}$, $265.33^{\circ}-274.68^{\circ}$; and $\beta\sim 77.32^{\circ}-80.33^{\circ}$, $99.09.33^{\circ}-102.78^{\circ},257.35^{\circ}-260.74^{\circ}$, $279.14^{\circ}-282.72^{\circ}$. For a better visualisation of the allowed and forbidden regions of the predicted parameters, we refer to Fig. (1).
Though the mass parameters are taken as input, yet a graphical visualisation of the mass spectra is put forward in Fig. (2). In addition, we generate the correlation plots to visualize the predictions of the studied parameters (see Figs. (2)-(2)). We know that the Jarlskog invariant ($J_{CP}$) [26], that represents the amount of the leptonic CP violation can be obtained in terms of standard parametrization as shown below,
@@
$\theta_{23}/^{\circ}$ 42.17 47.61 42.19 47.47
$\delta/^{\circ}$ 5.43 354.52 5.67 353.83
$\alpha/^{\circ}$ 85.37 274.68 85.53 274.56
$\beta/^{\circ}$ 77.32 282.72 77.36 282.59
$m_{\beta\beta}/\text{meV}$ 16.10 28.43 16.16 28.21
@@ 3.1 Effective Majorana Neutrino Mass
The neutrino-less double beta decay ($0\nu\beta\beta$)[27], is a lepton number violating decay. Its discovery would justify the Majorana nature of neutrinos. The decay rate $\Gamma$ is determined by the phase space factor $G^{0\nu}$, the nuclear matrix element $M^{0\nu}$, and the effective Majorana mass $m_{\beta\beta}$, following the relation: $\Gamma\sim G^{0\nu}.M^{0\nu}.m^{2}_{\beta\beta}$.
@@
where, $m_{1}$, $m_{2}$ and $m_{3}$ are the three mass eigenvalues. The $U_{11}$, $U_{12}$ and $U_{13}$ are the elements of the PMNS matrix that contains the information of $\alpha$ and $\beta$. Several experiments have provided the upper bounds of $m_{\beta\beta}$: SuperNEMO(Se82) as $67-131$ meV, GERDA(Ge76) as $104-228$ meV, EXO-200(Xe136) as $111-477$ meV, CUORE(Te130) as $75-350$ meV and KamLAND-Zen(Xe136) as $61-165$ meV [28, 29, 30, 31, 32, 33, 34]. In this regard, we visualize the prediction of the parameter $m_{\beta\beta}$ from the proposed texture (see Fig. 2). It is important to mention that the prediction of $m_{\beta\beta}$ from the proposed texture $(16.10-28.43)$ meV partially falls within the sensitivity of future experiments such as LEGEND-1000 [35].
@@ 3.2 CP Asymmetry Parameter
The transition probability $P(\nu_{\mu}\rightarrow\nu_{e})$ can be can be expressed in terms of the physical parameters [36] as shown below,
@@
The parameter $a=G_{F}N_{e}/\sqrt{2}$ depends on the medium through which the neutrino propagates and arises due to matter effects in neutrino propagation through the Earth. The factor $G_{F}$ stands for the Fermi constant and $N_{e}$ represents the electron number density of the medium. For the earth, $a$ is approximately $3500{\rm km}^{-1}$[36]. The CP asymmetry parameter can be expressed in terms of the physical parameters as shown below,
@@
Figure 4: Shows the correlation plots between (a) $|A|$ vs $r$. (b) $|G|$ vs $|F|$. (c) $Arg[A]$ vs $\theta$. (d) $Arg[F]$ vs $\theta$. (e) $Arg[G]$ vs $\theta$. The red data points signify texture prediction (TP) while the blue ones signify the corrections from model sector(MP).
@@ 4.1 The Yukawa Lagrangian
We start with a a framework where the SM gauge group is extended to $SU(2)_{L}\times U(1)_{Y}\times A_{4}\times Z_{10}$. In this regard, we extend the field content of the SM by introducing right-handed neutrinos and extra scalar fields. The transformation properties of the extended field content are summarised in Table (2). The product rules under $A_{4}$ group can be found in A. The desired Yukawa Lagrangian borne out of the said framework assumes the following form,
@@
$(\bar{D}_{l_{L}}\tilde{H})\nu_{\mu_{R}}\chi+\,\frac{y_{3}}{\Lambda}(\bar{D}_{l_{L}}\tilde{H})\nu_{\tau_{R}}\psi+\frac{y_{a}}{2}(\overline{\nu^{c}}_{e_{R}}\nu_{e_{R}})\xi+\frac{y_{b}}{2}[(\overline{\nu^{c}}_{\mu_{R}}\nu_{\tau_{R}})$
$+(\overline{\nu^{c}}_{\tau_{R}}\nu_{\mu_{R}})]\xi\kappa+\frac{y_{c}}{2}(\overline{\nu^{c}}_{\mu_{R}}\nu_{\mu_{R}})\eta+y_{T_{2}}(\bar{D}_{l_{L}}D_{l_{L}}^{c})\,i\tau^{2}\Delta+h.c.,$
where, $\Lambda$ is the cut-off scale of the theory. The Lagrangian is constructed in leading order. The presence of $Z_{10}$ [37] in the model restricts the undesirable terms that are permitted by $A_{4}$. Keeping in mind the targeted neutrino mass matrix in Eq. (1), the vacuum expectation values (vev) for the scalar fields are chosen as: $\langle H\rangle_{0}=v_{H}(1,1,1)^{T}$, $\langle\Delta\rangle_{0}=v_{\Delta}(v_{1},v_{2},0)^{T}$, $\langle\chi\rangle_{0}=v_{\chi}$, $\langle\psi\rangle_{0}=v_{\psi}$, $\langle\xi\rangle_{0}=v_{\xi}$, $\langle\eta\rangle_{0}=v_{\eta}$ and $\langle\kappa\rangle_{0}=v_{\kappa}$ respectively. The potential that shelters the chosen vacuum alignments is discussed in subsection 4.3.
@@
$Z_{10}$ 0 $(0,0,0)$ $(0,4,3)$ 00 0 6 $7$ 00 $2$ 3
Table 2: The transformation properties of various fields under $SU(2)_{L}\times U(1)_{Y}\times A_{4}\times Z_{10}$ group.
We derive the charged lepton mass matrix($M_{l}$) as shown in the following,
@@
$U_{l_{L}}=\frac{1}{\sqrt{3}}\begin{bmatrix}1&1&1\\ 1&\omega&\omega^{2}\\ 1&\omega^{2}&\omega\\ \end{bmatrix},\quad U_{l_{R}}=diag\begin{bmatrix}1,&1,&1\end{bmatrix}.$
The Dirac neutrino mass matrix($M_{D}$) and the right-handed neutrino mass matrix($M_{R}$) are obtained as follows,
$M_{D}=\begin{bmatrix}y_{1}v_{H}&\frac{y_{2}}{\Lambda}v_{H}v_{\chi}&\frac{y_{3}}{\Lambda}v_{H}v_{\psi}\\ y_{1}v_{H}&\frac{y_{2}}{\Lambda}\omega v_{H}v_{\chi}&\frac{y_{3}}{\Lambda}\omega^{2}v_{H}v_{\psi}\\ y_{1}v_{H}&\frac{y_{2}}{\Lambda}\omega^{2}v_{H}v_{\chi}&\frac{y_{3}}{\Lambda}\omega v_{H}v_{\psi}\\ \end{bmatrix},\quad M_{R}=\begin{bmatrix}\frac{y_{a}}{2}v_{\xi}&0&0\\ 0&\frac{y_{c}}{2}v_{\eta}&\frac{y_{b}}{2}v_{\xi}v_{\kappa}\\ 0&\frac{y_{b}}{2}v_{\xi}v_{\kappa}&0\\ \end{bmatrix}.$
@@
$M_{\nu s}=-M_{D}M_{R}^{-1}M_{D}^{T}\,\,\,+\,\,\begin{bmatrix}0&0&y_{T_{2}}v_{1}\\ 0&0&y_{T_{2}}v_{2}\\ y_{T_{2}}v_{1}&y_{T_{2}}v_{2}&0\\ \end{bmatrix}.$
@@
$\scriptstyle M_{\nu}=\begin{bmatrix}-\frac{6v^{2}y_{1}^{2}}{y_{a}v_{\xi}}+\frac{2v_{1}y_{T}}{3}+\frac{2v_{2}y_{T}}{3}&\frac{y_{T}}{3}\sqrt{v_{1}^{2}-v_{2}v_{1}+v_{2}^{2}}\,e^{i\phi}&\frac{y_{T}}{3}\sqrt{v_{1}^{2}-v_{2}v_{1}+v_{2}^{2}}\,e^{-i\phi}\\ \frac{y_{T}}{3}\sqrt{v_{1}^{2}-v_{2}v_{1}+v_{2}^{2}}\,e^{i\phi}&\frac{6v^{2}y_{3}^{2}y_{c}v_{\eta}v_{\psi}^{2}}{\Lambda^{2}y_{b}^{2}v_{\kappa}^{2}v_{\xi}^{2}}+y_{T}(\frac{2v_{2}}{3}-\frac{v_{1}}{3}-\frac{iv_{1}}{\sqrt{3}})&-\frac{6v^{2}y_{2}y_{3}v_{\chi}v_{\psi}}{\Lambda^{2}y_{b}v_{\kappa}v_{\xi}}-\frac{y_{T}}{3}(v_{1}+v_{2})\\ \frac{y_{T}}{3}\sqrt{v_{1}^{2}-v_{2}v_{1}+v_{2}^{2}}\,e^{-i\phi}&-\frac{6v^{2}y_{2}y_{3}v_{\chi}v_{\psi}}{\Lambda^{2}y_{b}v_{\kappa}v_{\xi}}-\frac{v_{1}y_{T}}{3}-\frac{v_{2}y_{T}}{3}&\frac{2y_{T}}{3}\sqrt{v_{1}^{2}-v_{2}v_{1}+v_{2}^{2}}\,e^{i\phi}\end{bmatrix},$ (3)
where, $\phi=-\tan^{-1}\left[\sqrt{3}v_{1}/(v_{1}-2v_{2})\right]$. The above model enhanced mass matrix carrying sixteen model parameters involving eight vevs, seven Yukawa couplings and the cut-off scale, resembles the exact form of the neutrino mass matrix texture expressed in terms of $r,\theta,A,F$ and $G$ as presented in Eq. (1). It’s worth mentioning that the independence of the texture parameters are maintained while expressing them in terms of the model parameters.
The given model incorporates several scalar fields, which play a crucial role in maintaining the independence of the texture parameters. In addition, to cut short certain terms from the right-handed neutrino mass matrix $(M_{R})$, we introduce some extra scalar fields in our model. This step is necessary to enhance the efficiency of the model. In the next section, we shall try to relate the parameters from texture and model sector.

### v3

@@ Abstract
Under the exponential parametrization scheme $m^{\nu}_{ij}\sim re^{i\theta}$, we propose a unique neutrino mass matrix texture that exhibits four correlations among its elements. The mixing scheme obtained from the proposed texture is consistent with experimental observations. We construct a concrete model based on the $SU(2)_{L}\times U(1)_{Y}\times A_{4}\times Z_{10}\times Z_{7}$ group within the framework of the seesaw mechanism.
@@ 1 Introduction
The neutrino oscillation suggests that neutrinos change their identity as they traverse space Pontecorvo:1957cp . This phenomenon can be understood from the quantum mechanics framework which tells us that the three flavour states of neutrinos ($\nu_{l=e,\mu,\tau}$) are the mixture of three mass eigenstates $(\nu_{i=1,2,3})$ having definite masses ($m_{i=1,2,3}$). The neutrino oscillation predicts the existence of non-zero mass, which goes against the predictions of the Standard Model (SM) Glashow:1961tr ; Weinberg:1967tq ; Salam:1968rm . In this regard, we need to go beyond the SM framework to understand the non-zero neutrino masses. The neutrino mass matrix ($M_{\nu}$) originates from the Yukawa Lagrangian, which can be understood from the see-saw mechanism Yoshimura:1978ex ; Akhmedov:1999tm ; Schechter:1980gr ; Schechter:1981cv ; Minkowski:1977sc ; Gell-Mann:1979vob ; Mohapatra:1979ia ; Magg:1980ut ; Lazarides:1980nt . In general, a Majorana neutrino mass matrix contains the information of twelve parameters: three mass eigenvalues ($m_{i=1,2,3}$), three mixing angles ($\theta_{12},\theta_{13},\theta_{23}$), three CP phases ($\delta,\alpha,\beta$), and three unphysical phases ($\phi_{1},\phi_{2},\phi_{3}$). Recent experiments predict three mixing angles and two mass-squared differences ($\Delta\,m_{21}^{2}$ and $\Delta\,m_{31}^{2}$) precisely Esteban:2024eli . However, the octant of $\theta_{23}$, neutrino mass ordering and a precise bound of $\delta$ are yet to be determined. In addition, the oscillation experiments do not measure three individual mass eigenvalues and the Majorana phases.
@@
However, the $M_{\nu}$ carries the information of all nine physical parameters, and by imposing certain constraints on $M_{\nu}$, some of the physical parameters can be predicted. Various constraints, such as texture zeroes Hagedorn:2005kz ; Grimus:2004hf ; Xing:2002ta , vanishing minors Lashin:2009yd ; Dev:2011hf , hybrid textures Liu:2013oxa , and $\mu-\tau$ symmetry Harrison:2002er are widely studied in the literature. The experiments rule out the $\mu-\tau$ symmetry due to its prediction of $\theta_{13}$ being zero. In this regard, there are several attempts made to explore deviations from the exact $\mu-\tau$ symmetry Dey:2022qpu ; Chakraborty:2023msb ; Chakraborty:2024rgt . In the present work, we explore different types of constraints, deviating from the existing literature. We adhere to the exponential parametrization of the neutrino mass matrix $M_{\nu_{ij}}=r_{ij}e^{i\theta_{ij}}$, where, $r_{ij}$ represents the absolute value and $\theta$ stands for argument of the matrix elements. In contrast, a complex number $z$ can also be expressed as $z=x+iy$. In principle, physics is independent of parametrization. The underlying principles and phenomena do not alter regardless of the parametrization we choose. However, a suitable parametrization may lead to understand the underlying physics in a better way. Like several other constraints studied in the literature, the said parametrization scheme offers new insights into how physical parameters can be correlated in a different way. The idea of the said parametrization was first discussed in a model-independent way in Ref Chakraborty:2022ess . The latter discusses several promising textures from different categories, highlighting distinct correlations including the exponential parametrization. All the proposed textures are experimentally viable. However, we have not provided any concrete model in support of the said textures.
@@ 2.1 The Proposed Texture
To draw information on the physical parameters, we need to diagonalize $M:M^{d}=V^{T}.M.V$, where, $V$ is a $\times 3$ 3 unitary matrix known as Pontecorvo-Maki-Nakagawa-Sakata (PMNS) matrix Maki:1962mu . In general, the PMNS matrix is defined in flavour basis i.e., the basis where the charged lepton mass matrix is diagonal. The Particle Data Group (PDG) ParticleDataGroup:2020ssz has adopted a parametrization for $V$, where, $V$ is parametrized using three angles and six phases as shown below,
@@ 2.2 Numerical Analysis
The four correlations appearing in $M_{\nu}$ of Eq. (1), are manifested as four transcendental equations involving nine physical parameters: $f_{i}(\theta_{12},\theta_{13},\theta_{23},m_{1},m_{2},m_{3},\delta,\alpha,\beta)=0$, where, $i=1,2.3,4$. We solve these equations by the Newton-Raphson method to generate a sufficient number of random numbers for the parameters $\alpha$, $\beta$, $\delta$ and $\theta_{23}$. For this, we set the mixing angles $\theta_{12}$ and $\theta_{13}$ at their respective $3\sigma$ bound. The mass eigenvalues $m_{1}$, $m_{2}$ and $m_{3}$ are estimated from two mass squared differences $\Delta m_{21}^{2}$ and $\Delta m_{31}^{2}$, such that $\Sigma m_{i}$ is consistent with the cosmological data Planck:2018vyg .
The analysis leads to some interesting predictions of the observable parameters. We see that $\theta_{23}$ lies within $42^{\circ}$ to $47^{\circ}$ approximately (within the $3\sigma$ range Esteban:2024eli ). The $\delta$ experiences two narrow bounds: $5.43^{\circ}$-$12.32^{\circ}$; and $347.52^{\circ}$-$354.52^{\circ}$. Thus, we see that within the allowed $3\sigma$ bound for $\delta$ given by the experiments, the above texture defines a forbidden region also. Similarly, we see that the Majorana phases $\alpha$ and $\beta$ have got two and four allowed bounds respectively. We see, $\alpha\sim 85.37^{\circ}-94.66^{\circ}$, $265.33^{\circ}-274.68^{\circ}$; and $\beta\sim 77.32^{\circ}-80.33^{\circ}$, $99.09.33^{\circ}-102.78^{\circ},257.35^{\circ}-260.74^{\circ}$, $279.14^{\circ}-282.72^{\circ}$. For a better visualisation of the allowed and forbidden regions of the predicted parameters, we refer to Fig. (1).
Though the mass parameters are taken as input, yet a graphical visualisation of the mass spectra is put forward in Fig. (2). In addition, we generate the correlation plots to visualize the predictions of the studied parameters (see Figs. (2)-(2)). We know that the Jarlskog invariant ($J_{CP}$) Jarlskog:1985ht , that represents the amount of the leptonic CP violation can be obtained in terms of standard parametrization as shown below,
@@
$\theta_{23}/^{\circ}$ 42.17 47.61 42.39 47.48
$\delta/^{\circ}$ 5.43 354.52 5.91 354.29
$\alpha/^{\circ}$ 85.37 274.68 85.37 274.48
$\beta/^{\circ}$ 77.32 282.72 77.31 282.54
$m_{\beta\beta}/\text{meV}$ 16.10 28.43 16.34 28.35
@@ 3.1 Effective Majorana Neutrino Mass
The neutrino-less double beta decay ($0\nu\beta\beta$)Schechter:1981bd , is a lepton number violating decay. Its discovery would justify the Majorana nature of neutrinos. The decay rate $\Gamma$ is determined by the phase space factor $G^{0\nu}$, the nuclear matrix element $M^{0\nu}$, and the effective Majorana mass $m_{\beta\beta}$, following the relation: $\Gamma\sim G^{0\nu}.M^{0\nu}.m^{2}_{\beta\beta}$.
@@
where, $m_{1}$, $m_{2}$ and $m_{3}$ are the three mass eigenvalues. The $U_{11}$, $U_{12}$ and $U_{13}$ are the elements of the PMNS matrix that contains the information of $\alpha$ and $\beta$. Several experiments have provided the upper bounds of $m_{\beta\beta}$: SuperNEMO(Se82) as 6713167-13167 - 131 meV, GERDA(Ge76) as 104228104-228104 - 228 meV, EXO-200(Xe136) as $1$ 477 meV, CUORE(Te130) as 7535075-35075 - 350 meV and KamLAND-Zen(Xe136) as 6116561-16561 - 165 meV Ejiri:2020xmm ; Agostini:2022zub ; CUORE:2019yfd ; GERDA:2019ivs ; KamLAND-Zen:2016pfg ; SuperNEMO:2021hqx ; CUORE:2018ncg . In this regard, we visualize the prediction of the parameter $m_{\beta\beta}$ from the proposed texture (see Fig. 2). It is important to mention that the prediction of $m_{\beta\beta}$ from the proposed texture (16.1028.43)(16.10-28.43)( 16.10 - 28.43 ) meV partially falls within the sensitivity of future experiments such as LEGEND-1000 LEGEND:2021bnm .
@@ 3.2 CP Asymmetry Parameter
The transition probability $P(\nu_{\mu}\rightarrow\nu_{e})$ can be can be expressed in terms of the physical parameters Sinha:2018xof as shown below,
@@
The parameter $a=G_{F}N_{e}/\sqrt{2}$ depends on the medium through which the neutrino propagates and arises due to matter effects in neutrino propagation through the Earth. The factor $G_{F}$ stands for the Fermi constant and $N_{e}$ represents the electron number density of the medium. For the earth, $a$ is approximately $3500{\rm km}^{-1}$inha:2018xof . The CP asymmetry parameter can be expressed in terms of the physical parameters as shown below,
@@
[…]
@@ 4.1 The Yukawa Lagrangian
We start with a a framework where the SM gauge group is extended to $SU(2)_{L}\times U(1)_{Y}\times A_{4}\times Z_{10}\times Z_{7}$. In this regard, we extend the field content of the SM by introducing right-handed neutrinos and extra scalar fields. The transformation properties of the extended field content are summarised in Table (2). The product rules under $A_{4}$ group can be found in A. The desired Yukawa Lagrangian borne out of the said framework assumes the following form,
@@
$(\bar{D}_{l_{L}}\tilde{H})\nu_{\mu_{R}}\chi+\,\frac{y_{3}}{\Lambda}(\bar{D}_{l_{L}}\tilde{H})\nu_{\tau_{R}}\psi+\frac{y_{a}}{2}(\overline{\nu^{c}}_{e_{R}}\nu_{e_{R}})\xi+\frac{y_{b}}{2\Lambda}[(\overline{\nu^{c}}_{\mu_{R}}\nu_{\tau_{R}})$
$+(\overline{\nu^{c}}_{\tau_{R}}\nu_{\mu_{R}})]\xi\kappa+\frac{y_{c}}{2}(\overline{\nu^{c}}_{\mu_{R}}\nu_{\mu_{R}})\eta+y_{T_{2}}(\bar{D}_{l_{L}}D_{l_{L}}^{c})\,i\sigma^{2}\Delta+h.c.,$
where, $\Lambda$ is the cut-off scale of the theory. The Lagrangian is constructed in leading order. The presence of $Z_{10}$ Dey:2023bfa in the model restricts the undesirable terms that are permitted by $A_{4}$. Keeping in mind the targeted neutrino mass matrix in Eq. (1), the vacuum expectation values (vev) for the scalar fields are chosen as: $\langle H\rangle_{0}=v_{H}(1,1,1)^{T}$, $\langle\Delta\rangle_{0}=v_{\Delta}(v_{1},v_{2},0)^{T}$, $\langle\chi\rangle_{0}=v_{\chi}$, $\langle\psi\rangle_{0}=v_{\psi}$, $\langle\xi\rangle_{0}=v_{\xi}$, $\langle\eta\rangle_{0}=v_{\eta}$ and $\langle\kappa\rangle_{0}=v_{\kappa}$ respectively. The potential that shelters the chosen vacuum alignments is discussed in subsection 4.3.
@@
$Z_{10}$ 0 (0,0,0)(0,0,0)( 0 , 0 , 0 ) (0,4,3)(0,4,3)( 0 , 4 , 3 ) 0 0 6 777 0 222 3
$Z_{7}$2 (4,4,4)(4,4,4)( 4 , 4 , 4 ) (6,6,6)(6,6,6)( 6 , 6 , 6 ) 1 3 0 0 2 2 0
Table 2: The transformation properties of various fields under $SU(2)_{L}\times U(1)_{Y}\times A_{4}\times Z_{10}\times Z_{7}$ group.
[…]
@@
$U_{l_{L}}=\frac{1}{\sqrt{3}}\begin{bmatrix}e^{i\rho}&1&1\\ e^{i\rho}&\omega&\omega^{2}\\ e^{i\rho}&\omega^{2}&\omega\\ \end{bmatrix},\quad U_{l_{R}}=diag\begin{bmatrix}e^{i\rho},&1,&1\end{bmatrix}.$
Here, we include a phase $\rho$ in $U_{l_{L}}$ and $U_{l_{R}}$, motivated by the fact that the choices of the eigenvectors are not unique Dey:2022qpu ; Chakraborty:2023msb ; Chakraborty:2024rgt ; Goswami:2025jde . The presence of this arbitrary phase does not alter the original structure of $M_{l}$ rather specifies a preferred charged lepton basis. The presence of the phase in $U_{l_{L}}$ will certainly affect the neutrino mass matrix in flavour basis, i.e., the basis where the $M_{l}$ is diagonal. Thus, a proper choice of the phase will help us to obtain the desired neutrino mass matrix. In the present work, we set $\rho=\pi$.
[…]
$M_{D}=\begin{bmatrix}y_{1}v_{H}&\frac{y_{2}}{\Lambda}v_{H}v_{\chi}&\frac{y_{3}}{\Lambda}v_{H}v_{\psi}\\ y_{1}v_{H}&\frac{y_{2}}{\Lambda}\omega v_{H}v_{\chi}&\frac{y_{3}}{\Lambda}\omega^{2}v_{H}v_{\psi}\\ y_{1}v_{H}&\frac{y_{2}}{\Lambda}\omega^{2}v_{H}v_{\chi}&\frac{y_{3}}{\Lambda}\omega v_{H}v_{\psi}\\ \end{bmatrix},\quad M_{R}=\begin{bmatrix}\frac{y_{a}}{2}v_{\xi}&0&0\\ 0&\frac{y_{c}}{2}v_{\eta}&\frac{y_{b}}{2\Lambda}v_{\xi}v_{\kappa}\\ 0&\frac{y_{b}}{2\Lambda}v_{\xi}v_{\kappa}&0\\ \end{bmatrix}.$
@@
$M_{\nu s}=-M_{D}M_{R}^{-1}M_{D}^{T}\,\,\,+\,\,\begin{bmatrix}0&0&y_{T_{2}}v_{2}\\ 0&0&y_{T_{2}}v_{1}\\ y_{T_{2}}v_{2}&y_{T_{2}}v_{1}&0\\ \end{bmatrix}.$
@@
$\scriptstyle M_{\nu}=\begin{bmatrix}-\frac{6v^{2}y_{1}^{2}}{y_{a}v_{\xi}}+\frac{2v_{1}y_{T}}{3}+\frac{2v_{2}y_{T}}{3}&\frac{y_{T}}{3}\sqrt{v_{1}^{2}-v_{2}v_{1}+v_{2}^{2}}\,e^{i\phi}&\frac{y_{T}}{3}\sqrt{v_{1}^{2}-v_{2}v_{1}+v_{2}^{2}}\,e^{-i\phi}\\ \frac{y_{T}}{3}\sqrt{v_{1}^{2}-v_{2}v_{1}+v_{2}^{2}}\,e^{i\phi}&\frac{6v^{2}y_{3}^{2}y_{c}v_{\eta}v_{\psi}^{2}}{\Lambda^{2}y_{b}^{2}v_{\kappa}^{2}v_{\xi}^{2}}+y_{T}(\frac{2v_{1}}{3}-\frac{v_{2}}{3}-\frac{iv_{2}}{\sqrt{3}})&-\frac{6v^{2}y_{2}y_{3}v_{\chi}v_{\psi}}{\Lambda^{2}y_{b}v_{\kappa}v_{\xi}}-\frac{y_{T}}{3}(v_{1}+v_{2})\\ \frac{y_{T}}{3}\sqrt{v_{1}^{2}-v_{2}v_{1}+v_{2}^{2}}\,e^{-i\phi}&-\frac{6v^{2}y_{2}y_{3}v_{\chi}v_{\psi}}{\Lambda^{2}y_{b}v_{\kappa}v_{\xi}}-\frac{y_{T}}{3}(v_{1}+v_{2})&\frac{2y_{T}}{3}\sqrt{v_{1}^{2}-v_{2}v_{1}+v_{2}^{2}}\,e^{i\phi}\end{bmatrix},$ (3)
where, $\phi=-\tan^{-1}\left[\sqrt{3}v_{2}/(v_{2}-2v_{1})\right]$. The above model enhanced mass matrix carrying sixteen model parameters involving eight vevs, seven Yukawa couplings and the cut-off scale, resembles the exact form of the neutrino mass matrix texture expressed in terms of $r,\theta,A,F$ and $G$ as presented in Eq. (1). It’s worth mentioning that the independence of the texture parameters are maintained while expressing them in terms of the model parameters.
In principle, the Weinberg like operator,
$\mathcal{L}_{W}=\frac{1}{\Lambda}\bar{D}_{l_{L}}\tilde{H}\tilde{H}^{T}D_{l_{L}}^{c},$
may appear in the model Vien:2025fiu . In order to restrict the said terms in our model, the $Z_{7}$ symmetry is invoked.
We wish to highlight that the present symmetry framework is designed in order to obtain the proposed texture of neutrino mass matrix in Eq. (1) in its exact form. To achieve this, both type-I and type-II seesaw mechanisms are incorporated. However, this enhances the field content. The texture parameters including $r$, $\theta$ and the other complex parameters $A$, $F$ and $G$ are all independent. The model parameters involving the Yukawa couplings and vevs etc. combine in specific ways and resemble the texture parameters. However, this step is challenging as this transformation might lead to the loss of independence of the texture parameters. To restrict this, the entry of additional scalar fields is inevitable. However, we don’t deny that another minimalistic approach to the problem is possible.
In the next section, we shall try to relate the parameters from texture and model sector.